SUMMER 2008 UNIVERSITY OF CHICAGO REU PROGRAM Apprentice 4 week program

CHECK ONE:

I will participate in the program; I understand that it is my responsibility to be present M-F 9:00-5:00 from June 23 to July 18, 2007 and that I will not engage in any other employment during that period without the explicit approval of the program director. I also understand that I am welcome to attend the program July 23 to August 15 and that a final short piece of mathematical writing (expository or research) is required.

I will not participate in the program

NAME:

ADDRESS:

E-MAIL: